# Efficient Hybrid Language Model Compression through Group-Aware SSM Pruning

### v1

@@ 3.1 Mamba Pruning
[…]
@@ 3. Group-Constrained Ranking: Within each Mamba group $\mathcal{G}_{g}$, sort heads by their scores:
8:     $\mathcal{R}_{g}^{\text{sel}}\leftarrow\text{first }k_{g}\text{ elements of }\mathcal{R}_{g}$
@@ 3.6 Accuracy Recovery with Knowledge Distillation (KD)
The output probability distribution of an LLM for a given token $x_{i}$ is computed as: $p(x_{i},\tau)=\frac{\exp\left(\frac{x_{i}}{\tau}\right)}{\sum_{j=1}^{|V|}\exp\left(\frac{x_{j}}{\tau}\right)}$, where $\tau$ is the softmax temperature and ${|V|}$ is the vocabulary size. Logit-based KD loss across the sequence of all output tokens is represented as: $L_{\text{logits}}=\frac{1}{L}\sum_{k=1}^{L}\text{FKLD}(p_{t}^{k}(x,\tau),p_{s}^{k}(x,\tau))$; here, $p_{t}^{k}(x,\tau)$ and $p_{s}^{k}(x,\tau)$ represent the teacher and student probability distributions on the $k^{th}$ token, respectively, and $L$ represents the sequence length.
@@ 4.4 Evaluation Summary
Tables 4 to 6 present accuracy comparisons between our compressed 4B hybrid model, other similar-sized community models, and the parent 8B hybrid model.
Our 4B model retains over 96% of the original 8B model’s accuracy while improving throughput by $\sim$1.4x. Compared to other similarly sized community models, it delivers state-of-the-art accuracy across knowledge, math, coding, commonsense reasoning, and reading comprehension tasks, despite being trained on up to $\sim$40x fewer tokens. It also achieves $\sim$2.2x higher throughput and $\sim$1.8x lower latency than the second-best Phi-4-4B model (Figures 1 and 8). The aligned version further leads in math, coding, instruction following, and tool-use tasks.

### v2

# Minitron-SSM: Efficient Hybrid Language Model Compression through Group-Aware SSM Pruning

@@ 3.1 Mamba Pruning
$\hat{x}$ $=\text{conv1d}(x)$ (10)
$\hat{B}$ $=\text{conv1d}(B)$ (11)
$\hat{C}$ $=\text{conv1d}(C)$ (12)
@@ 3. Group-Constrained Ranking: Within each Mamba group $\mathcal{G}_{g}$, sort heads by their scores:
[…]
@@ 3.6 Accuracy Recovery with Knowledge Distillation (KD)
The output probability distribution of an LLM for a given token $x_{i}$ is computed as: $p(x_{i},\tau)=\frac{\exp\left(\frac{x_{i}}{\tau}\right)}{\sum_{j=1}^{|V|}\exp\left(\frac{x_{j}}{\tau}\right)}$, where $\tau$ is the softmax temperature and |V|{|V|} is the vocabulary size. Logit-based KD loss across the sequence of all output tokens is represented as: $L_{\text{logits}}=\frac{1}{L}\sum_{k=1}^{L}\text{FKLD}(p_{t}^{k}(x,\tau),p_{s}^{k}(x,\tau))$; here, $p_{t}^{k}(x,\tau)$ and $p_{s}^{k}(x,\tau)$ represent the teacher and student probability distributions on the $k^{th}$ token, respectively, and $L$ represents the sequence length.
@@ 4.4 Evaluation Summary
Model Garak Score AEGIS Score
Nemotron-H-8B 70.75 99.83
Nemotron-H-4B 67.77 98.17
Table 7: Safety scores before and after compression.
Evaluation Summary. Tables 4 to 7 present accuracy comparisons between our compressed 4B hybrid model, other similar-sized community models, and the parent 8B hybrid model. As shown in Tables, our 4B model retains over 96% of the original 8B model’s accuracy, including safety scores on Garak and AEGIS, while improving throughput by $\sim$1.4x. Compared to other similarly sized community models, it delivers state-of-the-art accuracy across knowledge, math, coding, commonsense reasoning, and reading comprehension tasks, despite being trained on up to $\sim$40x fewer tokens. It also achieves $\sim$2.2x higher throughput and $\sim$1.8x lower latency than the second-best Phi-4-4B model (Figures 1 and 8). The aligned version further leads in math, coding, instruction following, and tool-use tasks.
@@
4.5 Generalizability to Mamba2
To evaluate the generalizability of our compression strategy to other models, we apply it to the Mamba2 1.3B model [3]. We prune the model to 780M parameters via SSM and embedding pruning, and then subsequently train the pruned model on 10.5B tokens. We compare our pruned 780M model to the Mamba2 780M and 1.3B models trained from scratch on 300B tokens [3].
As shown in Table 8, our compressed 780M model, despite being trained on significantly fewer tokens (10.5B vs. 300B), outperforms the 780M model trained from scratch and achieves an average score comparable to the original 1.3B model. These results provide further insights into the generalizability of our compression method.
Benchmark Mamba2 780M Mamba2 1.3B Compressed 780M
ARC Challenge 28.6 33.2 34.2
ARC Easy 54.7 60.6 60.4
CommonsenseQA 19.6 20.9 26.4
HellaSwag 54.7 59.9 50.4
OpenbookQA 36.4 37.0 34.2
PIQA 72.1 73.5 71.3
RACE 21.8 24.8 32.5
Social IQA 41.0 42.9 41.1
TruthfulQA MC2 38.1 36.1 38.9
Winogrande 58.0 60.1 57.7
Avg 42.5 44.9 44.7
Tokens 300B 300B 10.5B
Table 8: Comparison of our compressed Mamba2 780M model against Mamba2 780M and 1.3B models trained from scratch [3]. Despite being trained on significantly fewer tokens (10.5B vs. 300B), our compressed model achieves a better average score than the 780M baseline.